OpenStudy (anonymous):

What is the value of the expression 7–3? A. –21 B. Negative Exponents C. Negative Exponents D. 343

OpenStudy (anonymous):

@emmigrace222

OpenStudy (anonymous):

What is the value of the expression ? A. Negative Exponents B. Negative Exponents C. 25 D. 3125

OpenStudy (anonymous):

@emmigrace222

OpenStudy (anonymous):

@clara.university

OpenStudy (anonymous):

@cheergurl_99

OpenStudy (anonymous):

@CrazyCountryGirl

OpenStudy (anonymous):

umm hold on hun

OpenStudy (anonymous):

@emmigrace222

OpenStudy (anonymous):

what is the expression?

OpenStudy (anonymous):

1\5-5 _

OpenStudy (anonymous):

@Awesome781

OpenStudy (anonymous):

@secret66

OpenStudy (anonymous):

@okrzesik10

OpenStudy (anonymous):

Answer: = 1/5 − 5/1 = ((1 × 1) − (5 × 5)) / (5 × 1) = (1 - 25) / 5 = -24/5 = -24/5 = -4 4/5

OpenStudy (anonymous):

What is the value of the expression ? A. 27 B. 18 C. D. What is the value of the expression ? A. 27 B. 18 C. D.